Need help with a GPA question
If a student has a cumulative (overall) GPA of 2.6 and receives a 4.0 for first semester senior year - what does the cumulative GPA move to??
thanks
DDC

Answer by scott8148(3737)

if the course load is a constant - equal credit hours for ALL semesters - then there are 6 semesters @ 2.6 and 1 @ 4.0

let x="new GPA" ... 6(2.6)+1(4.0)=7(x)


if the credit hours are not constant (a more likely scenario) then put credit hours in place of semesters

(old credit hours)(2.6)+(new credit hours)(4.0)=(total credit hours)(x)